Opened 6 years ago

Last modified 6 years ago

#420 assigned flight processing

GB11/00, flight day 272b/2011, Little Rissington

Reported by: knpa Owned by: knpa
Priority: immediate Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors: knpa

Description

Data location: ~arsf/arsf_data/2011/flight_data/arsf_internal/GB11_00-2011_272b_Little_Rissington

Data arrived from ARSF via network transfer on 06/10/2011

GB11/00 Scientific objective: ARSF boresight flight.

GB11/00 Priority: urgent

PI: ARSF Internal

Notes: End of year boresight flight.

Sensors:

Camera (Flown but not required)
Eagle (Requested)
Hawk (Requested)
LiDAR (Requested)

Change History (9)

comment:1 Changed 6 years ago by knpa

Starting navigation processing.

comment:2 Changed 6 years ago by knpa

Navigation complete.

Created sol, camera sol, and sbet.

comment:3 Changed 6 years ago by knpa

Eagle 17 has a nonsense gps end point:

GPS Ending point = {209.578064, 113.273148}

comment:4 Changed 6 years ago by knpa

Starting HS boresight.

comment:5 Changed 6 years ago by emca

Processing LiDAR calibration

Initial comparison to ground control points shows an elevation offset of ~20cm

comment:6 Changed 6 years ago by emca

LiDAR Calibration:

Reprocessed navigation, improved solution but result not great. Reprocessed the LAS files and they still differs from ground control points, but less than before. Compared to ground points using lascontrol, not TScan. Will check on TScan when windows machine is free. Possibly elevation differences are due to navigation problems.

PRH values look okay - not great but reasonable.

comment:7 Changed 6 years ago by knpa

Processed hyperspectral lines with previous boresight and they look fine after sct correction.

e1 (South), e2 (West), e3 (East) checked. A couple of pixels movement at edges but no worse than result of previous boresight (a bit better in fact - possibly due to a bad DEM being used last time).

h6 (South), h7 (East), h8 (West) checked - same as eagle.

comment:8 Changed 6 years ago by emca

Due to the elevation differences the LiDAR calibration has been reflown on day 301 - see ticket #424

comment:9 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Milestone set to 2011 data processing completion
  • Owner changed from emca to knpa
  • Status changed from new to assigned

Have prepared project for archiving.

Note: See TracTickets for help on using tickets.